OpenStudy (anonymous):

Solve. (8/x)+(15/x^2)=-1

OpenStudy (anonymous):

x=-3, -5 :)

OpenStudy (anonymous):

add the left side by making the denominators the same. (multiply by one strategically) \[\frac{ 8 }{ x } * \frac{ x }{x } + \frac{ 15 }{ x^2 } = -1\] \[\frac{ 8x + 15 }{ x^2 } = -1\] \[x^2 + 8x + 15 = 0 = (x+5)(x+3)\]

OpenStudy (anonymous):